OpenStudy (anonymous):

Find n for a series for which a = 30, d = -4, and Sn = -210.

OpenStudy (anonymous):

\[Sn= n/2 [2a+d(n-1)]\]

OpenStudy (anonymous):

\[-420/n= 60-4n+4\]

OpenStudy (anonymous):

You'll get a quadratic equation: \[4n^2-64n-420=0\]